OpenStudy (davidusa):

How many different isosceles triangles have integer side lengths and a perimeter of 23?

OpenStudy (davidusa):

This is an AMC question from 2005.

OpenStudy (davidusa):

A) 2 B) 4 C) 6 D) 9 E) 11

OpenStudy (anonymous):

P = side+side+base 23 = side + side + base 23 = s + s + b 23 = 2s + b

OpenStudy (davidusa):

I thought, if it has to be isosceles, and have 23 perimeter, an odd number, the non equal side must be odd.

OpenStudy (davidusa):



OpenStudy (davidusa):

so x is 1,3,5, so on to 21.

OpenStudy (anonymous):

23 = 2s + b can be written as 23 = 2y + x if you like

OpenStudy (zarkon):

you cant go all the way to 21

OpenStudy (davidusa):

ohh!!! 2 sides sum has to more than other side sum. thanks

OpenStudy (zarkon):

right..the sum of any two sides is grater than the third side

OpenStudy (zarkon):

triangle inequality

OpenStudy (davidusa):

i got 11 triangles before.

OpenStudy (davidusa):

so now it is....

OpenStudy (davidusa):

6 triangles. am i right?

OpenStudy (zarkon):

yes

OpenStudy (davidusa):

thanks. so much.